Need help with 3C. Thank you!In a survey of 990 Canadian adults, 790 say that the energy situation in Canada is very or fairly serious. 1. Find the point
Answer #1

here p= 0.798 E = 0.03 Zal2 = 2.576 Now, Sample size = n = (Z*0 / E)?*p*(1-P) = (2.576/0.03)^2*0.798*(1-0.798) = 1188.51 = 11